40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 2 07:54:55 2010, steve, Update, MOPA, laser headtemp is up htemp.jpghtemp20d.jpg
    Reply  Wed Jun 2 11:28:31 2010, Koji, Update, MOPA, laser headtemp is up 
Message ID: 3035     Entry time: Wed Jun 2 11:28:31 2010     In reply to: 3033
Author: Koji 
Type: Update 
Category: MOPA 
Subject: laser headtemp is up 

Last night we stopped the air conditioning. It made HDTEMP increase.
Later we restored them and the temperature slowly recovered. I don't know why the recovery was so slow.

Quote:

Is the cooling line clogged? The chiller temp is 21C See 1 and 20 days plots

 

ELOG V3.1.3-